OpenStudy (anonymous):

0.56g of gas Q occupies 448cm^3 at standard temperature and pressure(STP).What is the relative molecular mass of Q?[molar volume of gas at STP=22.4dm^3mol^-1] A 14 B 28 C 32 D 64 @Abhisar @Somy

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@dan815

OpenStudy (somy):

\(\sf Volume = mole \times STP~ volume\) this formula is used when you need to find mole, or volume & u have the given condition, such as STP or RTP so rearranging the formula to make mole as subject \[\sf mole = \frac{ Volume }{ STP~ volume }\] you have volume of gas given convert it to dm3 by dividing it to 1000 u r given condition STP = 22.4 dm3 so now find mole

OpenStudy (somy):

you found the mole now use mole formula \[\sf mole = \frac{ mass }{ Molecular~mass }\] rearranging the formula to make molecular mass as subject \[\sf Molecular ~ mass = \frac{ mass }{ mole }\] you found the mole and you are given mass in the q thus use them and find Molecular mass

OpenStudy (anonymous):

\[\frac{ 448cm^3 }{ 22400cm^3 }==0.02mol\]

OpenStudy (anonymous):

\[\frac{ 0.56g }{ 0.02mol }=28gmol^{-1}\]

OpenStudy (anonymous):

Thnx @Somy

OpenStudy (somy):

no problem :)